------------------------------------------------------------ Date: Thu Jul 27 00:02:26 2000, Shift: 16:00-24:00 Supervisor: kh, with: tv, mh Log book#3 pages 93 - 97 Shift Summary: ------------------------------------------------------------ No beam during entire shift. TPC voltages off. Changed MTPC1 gating grid to 180 +- 60 ie set G13, G14 and G15 to 260, 180, 80 in M4 Changed ZDC trigger 4 to 50 mV from 100 mV. Unfortunately no beam to test effects. Installed TPC pedestals near end of shift. Installed later version of brat in operator account. Old version of brat was moved to brat_old and can be deleted as soon as we see that all is in order with newer version. Preliminary tests with files indicate it is in order. ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Thu Jul 27 2000 - 00:06:05 EDT